# Verification and Refinement of Natural Language Explanations through LLM-Symbolic Theorem Proving

### v2

@@ Abstract
Natural language explanations have become a proxy for evaluating explainable and multi-step Natural Language Inference (NLI) models. However, assessing the validity of explanations for NLI is challenging as it typically involves the crowd-sourcing of apposite datasets, a process that is time-consuming and prone to logical errors. To address existing limitations, this paper investigates the verification and refinement of natural language explanations through the integration of Large Language Models (LLMs) and Theorem Provers (TPs). Specifically, we present a neuro-symbolic framework, named Explanation-Refiner, that augments a TP with LLMs to generate and formalise explanatory sentences and suggest potential inference strategies for NLI. In turn, the TP is employed to provide formal guarantees on the logical validity of the explanations and to generate feedback for subsequent improvements. We demonstrate how Explanation-Refiner can be jointly used to evaluate explanatory reasoning, autoformalisation, and error correction mechanisms of state-of-the-art LLMs as well as to automatically enhance the quality of human-annotated explanations of variable complexity in different domains.
@@ 1 Introduction
A recent line of research in Natural Language Inference (NLI) focuses on developing models capable of generating natural language explanations in support of their predictions (Thayaparan et al., 2021; Chen et al., 2021; Valentino et al., 2022; Bostrom et al., 2022; Weir et al., 2023). Since natural language explanations can be used as a proxy to evaluate the underlying reasoning process of NLI models (Kumar and Talukdar, 2020; Zhao and Vydiswaran, 2021; Chen et al., 2021), researchers have proposed different methods for assessing their intrinsic quality (Wiegreffe and Marasovic, 2021; Camburu et al., 2020; Valentino et al., 2021; Atanasova et al., 2023; Quan et al., 2024; Dalal et al., 2024), including the adoption of language generation metrics for a direct comparison between models’ generated explanations and human-annotated explanations.
However, this process is subject to different types of limitations. First, the use of language generation metrics requires the crowd-sourcing of explanation corpora to augment existing NLI datasets Wiegreffe and Marasovic (2021), a process that is typically time-consuming and susceptible to errors (Liu et al., 2022; Zhao et al., 2023). Second, language generation metrics have been shown to fail capturing fine-grained properties that are fundamental for NLI such as logical reasoning, faithfulness, and robustness (Atanasova et al., 2023; Camburu et al., 2020; Chan et al., 2022; Quan et al., 2024). Third, human explanations in NLI datasets tend to be incomplete and contain logical errors that could heavily bias the evaluation Elazar et al. (2021); Valentino et al. (2021).
In this paper, we investigate the integration of state-of-the-art LLM-based explanation generation models for NLI with external logical solvers to jointly evaluate explanatory reasoning Pan et al. (2023a); Olausson et al. (2023); Jiang et al. (2024b) and enhance the quality of crowd-sourced explanations. In particular, we present a neuro-symbolic framework, named Explanation-Refiner, that augments a Theorem Prover (TP) with Large Language Models (LLMs) to investigate the following research questions: RQ1: “Can the integration of LLMs and TPs provide a mechanism for automatic verification and refinement of natural language explanations?”; RQ2: “Can the integration of LLMs and TPs improve the logical validity of human-annotated explanations?”; RQ3: “To what extent are state-of-the-art LLMs capable of explanatory reasoning, autoformalisation, and error correction for NLI in different domains?”. To answer these questions, Explanation-Refiner employs LLMs to generate and formalise explanatory sentences and to suggest potential inference strategies for building non-redundant, complete, and logically valid explanations for NLI. In turn, the TP is adopted to verify the validity of the explanations through the construction of deductive proofs and the generation of fine-grained feedback for LLMs.
@@
Our empirical analysis carried out on three NLI datasets of variable complexity (i.e., e-SNLI (Camburu et al., 2018), QASC (Khot et al., 2019), and WorldTree (Jansen et al., 2018)) reveals that external feedback from TPs is effective in improving the quality of natural language explanations, leading to an increase in logical validity using GPT-4 from 36% to 84%, 12% to 55%, and 2% to 37% (on e-SNLI, QASC, and WorldTree respectively). At the same time, the results demonstrate that integrating external TPs with LLMs can reduce errors in autoformalisation, with an average reduction of syntax errors of 68.67%, 62.31%, and 55.17%. Finally, we found notable differences in performance across LLMs and NLI datasets, with closed-sourced LLMs (i.e., GPT-4 and GPT-3.5) significantly outperforming open-source models (i.e., Mistral and LLama) on both explanatory reasoning and autoformalisation, along with a shared tendency of LLMs to struggle with increasing explanation complexity.
Figure 1: The overall pipeline of Explanation-Refiner: An NLI problem is converted into axioms and theorems as a theorem prover’s theory, along with some proof steps derived from a preliminary inference to send to a theorem prover. In case the proof fails (logically invalid), the erroneous step, along with the proof strategy and proof steps, are extracted as feedback to refine the explanatory sentences in a new iteration.
@@
    We utilise Neo-Davidsonian event semantics coupled with FOL to effectively translate natural language sentences into logical forms to minimises semantic information loss. Additionally, we introduce a novel method that leverages a theorem prover and a proof assistant for verifying NLI explanations and a syntactic refiner to minimise syntax errors in responses generated by LLMs.
@@
    We conduct a comprehensive series of experiments to assess Explanation-Refiner across five LLMs on three datasets, which include between 1 to 16 explanatory sentences. These experiments span a range of tasks from simple textual entailment to complex multiple-choice question answering in different context domains.
@@
    We perform extensive quantitative and qualitative analyses to explore the explanation refinement process. Our investigations delve into the LLMs’ inference capabilities, revealing the strengths and limitations of different models in producing verifiable and explainable logical reasoning for NLI.
@@ 2 Explanation Verification and Refinement
Explanation-based NLI is widely adopted to evaluate the reasoning process of multi-step inference models via the construction of natural language explanations. In this work, we refer to the following formalisation for Explanation-based NLI: given a premise sentence $p_{i}$, a hypothesis sentence $h_{i}$, and an explanation $E_{i}$ consisting of a set of facts $\{f_{1},f_{2},...,f_{n}\}$, the explanation $E_{i}$ is logically valid if and only if the entailment $p_{i}\cup E_{i}\models h_{i}$ holds. This entailment is considered verifiable if $\{p_{i},E_{i},h_{i}\}$ can be translated into a set of logical forms $\Phi$ that compose a theory $\Theta$. The validity of this theory, $\Theta$, is subsequently determined by a theorem prover, verifying whether $\Theta\vDash\psi$, where $\psi$ represents a logical consequence derived from the logical form of $h_{i}$.
We aim to automatically verify the logical validity of explanation $E_{i}$, and if $\Theta\vDash\psi$ is rejected by the theorem prover, a further refinement stage should be initiated to refine the facts $\{f_{1},f_{2},...,f_{n}\}$ based on external feedback, resulting in an updated explanation $E_{i}^{\prime}$. Thus, an explanation is accepted if all the facts are logically consistent, complementary and non-redundant to support the derivation.
@@ 3 Explanation-Refiner
To verify the logical validity and refine any logical errors in explanatory sentences for NLI tasks, we present a neuro-symbolic framework, Explanation-Refiner, to iteratively check and refine the explanation $E_{i}$ based on external feedback; Isabelle/HOL (Nipkow et al., 2002) proof assistant is applied as the external tool to make the logical deduction and provide feedback for explanation refinement.
Figure 1 shows an overview of our proposed framework. Given an NLI task, to evaluate the logical validity of the entailment, the LLM is prompted to perform an autoformalisation process that transforms natural language sentences into formal language represented in the form of Isabelle/HOL theory. Each fact $f\in E_{i}$ is converted into an axiom $a_{i}$, where each $a_{i}$ is an element of the set $A=\{a_{1},a_{2},...,a_{n}\}$. The premise $p_{i}$ and corresponding hypothesis $h_{i}$, is converted into a theorem for proving $p_{i}\wedge B\rightarrow h_{i}$, where $B\subseteq A$. A syntax refinement mechanism is subsequently applied to the previously transferred symbolic forms. The theorem prover is implemented as a checker to identify any syntax errors and provide these error details as feedback to an LLM, enabling the LLM to iteratively correct the syntax errors over a fixed number of iterations, denoted by $t$.
With the proof assistant, we can perform automated reasoning in the theorem prover by constructing proof steps. Step 3 initially generate a rough inference to state a preliminary proof strategy in natural language sentences and elicit the facts $f\in E_{i}$ which are non-redundant and essential for entailing the hypothesis $h_{i}$. Based on this preliminary proof strategy, the LLM is prompted to construct and formalise the proof steps for proving the theorem.
In step 5, the theorem prover will verify the constructed theory by attempting to prove the theorem. If it is solvable, we state it as a logically valid explanation. If the prover failed at one of the proof steps, we extract this failed step along with the applied axioms $B\subseteq A$ as the external feedback for an LLM. This feedback is used to refine the logical error within B and consequently refine the facts $f\in E_{i}$, which were previously converted from natural language.
@@
Figure 2: An example of representing the premise and hypothesis sentences in Isabelle/HOL theorem
@@
Figure 3: An example of proof constructed by the LLM for the Isabelle/HOL proof assistant to verify the hypothesis.
@@ 3.1 Autoformalisation
Preventing the loss of semantic information during the representation of natural language sentences in logical forms, such as FOL, poses significant challenges when using LLMs, particularly with long and complex sentences that are crucial for logical reasoning (Olausson et al., 2023). Neo-Davidsonian event semantics (Parsons, 1990) utilises event variables to represent the verb predicates and their corresponding object arguments as semantic roles. This approach establishes a predicate-argument structure that preserves the information content and faithfulness of complex sentences, closer to the surface form of the sentence. For example, the sentence ‘A wolf eating a sheep is an example of a predator hunting prey’ can be formalised as follows:
@@
In 1, the verbs are represented as the events ‘eating’ and ‘hunting,’ where the agent and patient arguments correspond to the entities performing and receiving the actions within these events, respectively. The logical form $\text{example}(e_{1},e_{2})$ explicitly captures the semantic meaning of this sentence: the event of a wolf eating a sheep as an exemplar of a predator hunting prey. Similarly, whenever there are no action verbs involved in a sentence, we utilise FOL to represent the static or descriptive aspects. For instance:
@@
The above logical forms correspond to the sentences ‘gravity is a kind of force’ and ‘greater means larger’, respectively.
@@
For the Isabelle/HOL theorem prover, a theory script is essential to facilitate the proof of a theorem. Therefore, we designate explanatory sentences as axioms:
@@
Additionally, as illustrated in Figure 2, both the premises and hypothesis are defined as constituting parts of the theorem. The ‘assumes asm’ clause generally comprises unquantified, specific propositions or conjunctions of propositions, which are recognised as known truths. The ‘show’ clause denotes the conclusion (hypothesis) for which we seek to establish proof through logical deductions based on the assumed propositions and axioms.
@@
Recent studies (Gou et al., 2024; Olausson et al., 2023) have revealed persistent syntax errors when prompting LLMs for code and symbolic form generation tasks. The method proposed by Pan et al. (2023a) utilises error messages from symbolic solvers to iteratively refine LLM outputs. Following a similar approach, we categorised the syntax errors into two distinct subdomains based on feedback from Isabelle: type unification errors and other syntax errors. Type unification errors primarily arise from mismatches between declared and actual argument types in logical clauses. Other syntax errors typically involve missing brackets, undefined entity names, or invalid logical symbols. Our process involves using Isabelle to identify syntax errors in the transferred theory, extracting these error messages, and then prompting the LLM with these messages along with few-shot examples. This guides the model on how to correct each type of syntax error over a series of iterations, allowing for continuous verification and refinement. Details of the autoformalisation prompts are described in Appendix A.4.1.
@@ 3.2 Proof Construction
A proof provides a detailed, step-by-step strategy that elucidates the logical connections and unification among axioms to support the reasoning process aimed at achieving the solver’s goal. Initially, we prompt the LLM to create a preliminary proof in natural language to assess how it infers the hypothesis and to identify which explanatory sentences are relevant, redundant, or unrelated. Based on this initial proof, we then guide the LLM to develop an Isabelle/HOL proof (figure 3) that utilise the Isabelle/HOL proof assistant to clearly demonstrate the explanatory sentences (axioms) required to prove the hypothesis. The general proof steps generated by an LLM are in the format ‘show $X$ using $Y$ by $Z$’, where the theorem prover is asked to prove $X$ given the assumptions $Y$, using the automated proof tactic $Z$. The proof tactic often applied is ‘blast’, which is one of Isabelle’s powerful FOL theorem proving tactics, enabling efficient and automated proof discovery across a range of logical forms (Paulson, 1999). Details of proof construction process prompts are described in Appendix A.4.2.
3.3 Verify and Refine
Finally, the constructed theory, which includes axioms, theorems, and proof steps, is submitted to the theorem prover for verification. If the theory is validated, it outputs a logically sound explanation. If the proof fails or timeouts, we extract the first error from the solver’s error message, identify the corresponding proof step, and locate the related explanatory sentences (axioms) from the theory. We begin by removing redundant and irrelevant facts that are not present in the preceding Isabelle/HOL proof steps or are declared as such in the text inference strategy. Then, we prompt the LLM to refine the explanatory sentences by providing it with the error message, the failed proof step, the associated proof strategy, and the relevant explanatory sentences for further iteration. This process is iterative and progressive; with each iteration, the framework addresses one or more logical errors, continually refining the explanatory sentences to ultimately yield a logically valid and verifiable explanation. Additional details on the prompts used for refinement are described in Appendix A.4.3.

### v4

@@ Abstract
Natural language explanations represent a proxy for evaluating explanation-based and multi-step Natural Language Inference (NLI) models. However, assessing the validity of explanations for NLI is challenging as it typically involves the crowd-sourcing of apposite datasets, a process that is time-consuming and prone to logical errors. To address existing limitations, this paper investigates the verification and refinement of natural language explanations through the integration of Large Language Models (LLMs) and Theorem Provers (TPs). Specifically, we present a neuro-symbolic framework, named Explanation-Refiner, that integrates TPs with LLMs to generate and formalise explanatory sentences and suggest potential inference strategies for NLI. In turn, the TP is employed to provide formal guarantees on the logical validity of the explanations and to generate feedback for subsequent improvements. We demonstrate how Explanation-Refiner can be jointly used to evaluate explanatory reasoning, autoformalisation, and error correction mechanisms of state-of-the-art LLMs as well as to automatically enhance the quality of explanations of variable complexity in different domains. [^1] Code and data are available at: https://github.com/neuro-symbolic-ai/explanation_refinement
@@ 1 Introduction
A recent line of research in Natural Language Inference (NLI) focuses on developing models capable of generating natural language explanations in support of their predictions (Thayaparan et al., 2021; Chen et al., 2021; Valentino et al., 2022a; Bostrom et al., 2022; Weir et al., 2023). Since natural language explanations can be used as a proxy to evaluate the underlying reasoning process of NLI models (Kumar and Talukdar, 2020; Zhao and Vydiswaran, 2021; Chen et al., 2021), researchers have proposed different methods for assessing their intrinsic quality (Camburu et al., 2020; Wiegreffe and Marasovic, 2021; Valentino et al., 2021; Atanasova et al., 2023; Quan et al., 2024; Dalal et al., 2024), including the adoption of language generation metrics for a direct comparison between models’ generated explanations and human-annotated explanations.
However, this process is subject to different types of limitations. First, the use of language generation metrics requires the crowd-sourcing of explanation corpora to augment existing NLI datasets Wiegreffe and Marasovic (2021), a process that is time-consuming and susceptible to errors (Valentino et al., 2021; Liu et al., 2022; Zhao et al., 2023). Second, language generation metrics have been shown to fail capturing fine-grained properties that are fundamental for NLI such as logical reasoning, faithfulness, and robustness (Camburu et al., 2020; Chan et al., 2022; Atanasova et al., 2023; Quan et al., 2024). Third, human explanations in NLI datasets tend to be incomplete and contain logical errors that could heavily bias the evaluation Elazar et al. (2021); Valentino et al. (2021).
In this paper, we investigate the integration of state-of-the-art LLM-based explanation generation models for NLI with external logical solvers to jointly evaluate explanatory reasoning Pan et al. (2023a); Olausson et al. (2023); Jiang et al. (2024b) and enhance the quality of crowd-sourced explanations. In particular, we present a neuro-symbolic framework, named Explanation-Refiner, that integrates a Theorem Prover (TP) with Large Language Models (LLMs) to investigate the following research questions: RQ1: “Can the integration of LLMs and TPs provide a mechanism for automatic verification and refinement of natural language explanations?”; RQ2: “Can the integration of LLMs and TPs improve the logical validity of human-annotated explanations?”; RQ3: “To what extent are state-of-the-art LLMs capable of explanatory reasoning, autoformalisation, and error correction for NLI in different domains?”. To answer these questions, Explanation-Refiner employs LLMs to generate and formalise explanatory sentences and to suggest potential inference strategies for building non-redundant, complete, and logically valid explanations for NLI. In turn, the TP is adopted to verify the validity of the explanations through the construction of deductive proofs and the generation of fine-grained feedback for LLMs.
@@
Our empirical analysis, carried out on three NLI datasets of variable complexity (i.e., e-SNLI (Camburu et al., 2018), QASC (Khot et al., 2019), and WorldTree (Jansen et al., 2018)), reveals that external feedback from TPs is effective in improving the quality of natural language explanations, leading to an increase in logical validity using GPT-4 from 36% to 84%, 12% to 55%, and 2% to 37% (on e-SNLI, QASC, and WorldTree respectively). At the same time, the results demonstrate that integrating external TPs with LLMs can reduce errors in autoformalisation, with an average reduction of syntax errors of 68.67%, 62.31%, and 55.17%. Finally, we found notable differences in performance across LLMs and NLI datasets, with closed-sourced LLMs (i.e., GPT-4 and GPT-3.5) significantly outperforming open-source models (i.e., Mistral and LLama) on both explanatory reasoning and autoformalisation, along with a shared tendency of LLMs to struggle with increasing explanation complexity.
Figure 1: The overall pipeline of Explanation-Refiner: An NLI problem is converted into axioms and theorems for a theorem prover, along with some proof steps derived from a preliminary inference. In case the proof fails (logically invalid), the erroneous steps along with the constructed proof strategy are used as feedback to refine the explanation in a new iteration.
@@
    We integrate Neo-Davidsonian event semantics coupled with FOL to effectively translate natural language sentences into logical forms to minimise semantic information loss. Additionally, we introduce a novel method that leverages a theorem prover and a proof assistant for verifying NLI explanations and a syntactic refiner to minimise syntax errors in responses generated by LLMs.
@@
    We conduct a comprehensive series of experiments with Explanation-Refiner across five LLMs and three datasets, including 1 to 16 explanatory sentences, covering tasks from textual entailment to complex multiple-choice question answering in various domains.
@@
    We perform extensive analyses to explore the explanation refinement process, characterising the LLMs’ inference capabilities and revealing the strengths and limitations of different models in producing verifiable, explainable logical reasoning for NLI.
@@ 2 Explanation Verification and Refinement
Explanation-based NLI is widely adopted to evaluate the reasoning process of multi-step inference models via the construction of natural language explanations. In this work, we refer to the following formalisation for Explanation-based NLI: given a premise sentence $p_{i}$, a hypothesis sentence $h_{i}$, and an explanation $E_{i}$ consisting of a set of facts $\{f_{1},f_{2},...,f_{n}\}$, the explanation $E_{i}$ is logically valid if and only if the entailment $p_{i}\cup E_{i}\models h_{i}$ holds. This entailment is considered verifiable if $\{p_{i},E_{i},h_{i}\}$ can be translated into a set of logical formulae $\Phi$ that compose a theory $\Theta$. The validity of the theory $\Theta$ is subsequently determined by a theorem prover, verifying whether $\Theta\vDash\psi$, where $\psi$ represents a logical consequence derived from the logical form of $h_{i}$.
In this paper, we aim to automatically verify the logical validity of an explanation $E_{i}$. To this end, if $\Theta\vDash\psi$ is rejected by the theorem prover, a further refinement stage should be initiated to refine the facts $\{f_{1},f_{2},...,f_{n}\}$ based on external feedback, resulting in an updated explanation $E_{i}^{\prime}$. Thus, an explanation is accepted if all the facts are logically consistent, complementary and non-redundant to support the derivation.
@@ 3 Explanation-Refiner
To verify the logical validity and refine any logical errors in explanatory sentences for NLI tasks, we present a neuro-symbolic framework that iteratively checks and refines the explanation $E_{i}$ based on external feedback. Figure 1 shows an overview of our proposed framework. Given an NLI task, to evaluate the logical validity of the entailment, the LLM is prompted to perform an autoformalisation process that transforms natural language sentences into formal language represented in the form of an Isabelle/HOL theory. Each fact $f\in E_{i}$ is converted into an axiom $a_{i}$, where each $a_{i}$ is an element of the set $A=\{a_{1},a_{2},...,a_{n}\}$. The premise $p_{i}$ and corresponding hypothesis $h_{i}$, is converted into a theorem for proving $p_{i}\wedge B\rightarrow h_{i}$, where $B\subseteq A$. A syntax refinement mechanism is subsequently applied to the previously transferred symbolic forms. The theorem prover is implemented as a checker to identify any syntax errors and provide these error details as feedback to an LLM, enabling the LLM to iteratively correct the syntax errors over a fixed number of iterations, denoted by $t$.
We can then perform automated reasoning via the theorem prover. To this end, in step 3 we use the LLM to generate a rough inference that states a preliminary proof strategy in natural language and elicit the facts $f\in E_{i}$ which are sufficient and necessary for entailing the hypothesis $h_{i}$. Based on this preliminary proof strategy, the LLM is prompted to construct and formalise the proof steps for proving the theorem. In step 5, the theorem prover will verify the constructed theory by attempting to prove the theorem. If it is solvable, we consider it a logically valid explanation. If the prover failed at one of the proof steps, we adopt the failed steps along with the applied axioms $B\subseteq A$ as an external feedback for the LLM. This feedback is used to refine the logical errors and consequently refine the facts $f\in E_{i}$.
@@
Figure 2: An example of representing the premise and hypothesis sentences in Isabelle/HOL theorem includes a proof constructed by the LLM for verifying the hypothesis.
@@ 3.1 Autoformalisation
Preventing the loss of semantic information during the representation of natural language sentences in logical forms, such as FOL, poses significant challenges when using LLMs, particularly with long and complex sentences that are crucial for logical reasoning (Olausson et al., 2023). Neo-Davidsonian event semantics (Parsons, 1990) focused on event variables to represent the verb predicates and their corresponding object arguments as semantic roles. This approach establishes a predicate-argument structure that preserves the information content and faithfulness of complex sentences, closer to the surface form of the sentence Quan et al. (2024). For example, the sentence ‘A wolf eating a sheep is an example of a predator hunting prey’ can be formalised as follows:
@@
In 1, the verbs are represented as the events ‘eating’ and ‘hunting,’ where the agent and patient arguments correspond to the entities performing and receiving the actions within these events, respectively. The logical form $\text{example}(e_{1},e_{2})$ explicitly captures the semantic meaning of this sentence: the event of a wolf eating a sheep as an exemplar of a predator hunting prey. Similarly, whenever there are no action verbs involved in a sentence, we use FOL to represent the static or descriptive aspects. For instance:
@@
The above logical forms correspond to the sentences ‘gravity is a kind of force’ and ‘greater means larger’.
@@
A theory script for the Isabelle/HOL theorem prover contains theorems that need to be proven from some axioms. Therefore, we adopt the sentences in an explanation to construct the set of axioms. For instance:
@@
In addition, as illustrated in Figure 2, both the premise and the hypothesis constitute parts of the theorem to be proven. In particular, the ‘assumes asm’ clause includes unquantified, specific propositions or conjunctions of propositions which are recognised as known truths (i.e., premises). On the other hand, the ‘show’ clause denotes the conclusion (i.e., hypothesis) for which we seek to build a proof through logical deductions based on the assumed propositions and axioms.
@@
Recent studies (Olausson et al., 2023; Gou et al., 2024) have revealed persistent syntax errors when prompting LLMs for code and symbolic form generation tasks. We categorised the syntax errors into two distinct subdomains based on feedback from Isabelle: type unification errors and other syntax errors. Type unification errors primarily arise from mismatches between declared and actual argument types in logical clauses. Other syntax errors typically involve missing brackets, undefined entity names, or invalid logical symbols. Our process involves using Isabelle to identify syntax errors in the transferred theory, extracting these error messages, and then prompting the LLM with these messages along with few-shot examples. This guides the model on how to correct each type of syntax error over a series of iterations, allowing for continuous verification and refinement. Details of the autoformalisation prompts are described in Appendix A.4.1.
@@ 3.2 Proof Construction
A proof provides a detailed step-by-step strategy that elucidates the logical connections and unification among axioms to support the reasoning process aimed at achieving the solver’s goal. Initially, we prompt the LLM to create a preliminary proof in natural language to assess how it infers the hypothesis and to identify which explanatory sentences are relevant, redundant, or unrelated. Based on this initial inference, we then guide the LLM to develop a formal proof (Figure 2) that integrates with Isabelle/HOL to verify the explanatory sentences (axioms) that are required to derive the hypothesis. The general proof steps generated by an LLM are in the format ’show $X$ using $Y$ by $Z$’, where the theorem prover is asked to prove $X$ given the assumptions $Y$, using the automated proof tactic $Z$. The proof tactic often applied is ’blast’, which is one of broader Isabelle’s FOL theorem proving tactics(Paulson, 1999). Additional details of the proof construction process and the prompts used to guide the LLMs are described in Appendix A.4.2.
3.3 Verification and Refinement
Finally, the constructed theory, which includes axioms, theorems, and proof steps, is submitted to the theorem prover for verification. If the theory is validated, it outputs a logically valid explanation. If the proof fails or timeouts, we extract the first error from the solver’s error message, identify the corresponding proof step, and locate the related explanatory sentences (axioms) from the theory. We begin by removing redundant and irrelevant facts that are not present in the preceding Isabelle/HOL proof steps or are declared as such in the text inference strategy. Then, we prompt the LLM to refine the explanatory sentences by providing it with the error message, the failed proof step, the associated proof strategy, and the relevant explanatory sentences for further iteration. This process is iterative and progressive; with each iteration, the framework addresses one or more logical errors, continually refining the explanatory sentences to ultimately yield a logically valid and verifiable explanation. Additional details on the prompts used for refinement are described in Appendix A.4.3.